OpenStudy (anonymous):

2/9 ÷ z/3 2/3z 2z/27 2/27 2/3

OpenStudy (solomonzelman):

Does you equation look like this: \(\large\color{black}{ \displaystyle \frac{2}{9}\div \frac{z}{3} }\) ?

OpenStudy (solomonzelman):

Closed Question? I would have helped, but oh well, I guess not :(